A 100 g aluminum calorimeter contains 250 g of water. The two substances are in thermal equilibrium at 10°C. Two metallic blocks are placed in the water. One is a 50 g piece of copper at 76°C. The other sample has a mass of 73 g and is originally at a temperature of 100°C. The entire system stabilizes at a final temperature of 20°C. Determine the specific heat of the unknown second sample.

J/kg·°C
Answer #1

mAl = 100 g, mw = 250 g, mcu = 50 g,m = 73 g,

specific heat capacity of water = 4186 J/kg.oC
specific heat capacity of Al = 900 J/kg.oC
specific heat capacity of copper = 387 J/kg.oC
Energy gain by system = energy lost by system

mA1Cp*(T2-T1) +mw*Cp(T2-T1) = -mcu*Cp*(T2-T1) - m*C*(T2-T1)
100*900*(20-10)+250*4186*(20-10) = -50*387*(20-76) -73*C*(20-100)
specific heat capacity of unkonwn material C = 1760 J/kg.oC